x^2 - 16x + 49 = 17 - x^2
Answer by unlockmath(1688)

x^2 -16x + 49 = 17 - x^2
2x^2 -16x + 49 = 17
2x^2 -16x + 32 = 0
2(x^2 -8x + 16) = 0
2 (x-4)(x-4) = 0
x = 4